OpenStudy (anonymous):

Okay, I need someone to set me in the right direction on this question... Find the number of equivalent radians in exact value: 270°

OpenStudy (johnweldon1993):

To convert from degrees to radians you do \[270 \times \frac{ \pi }{ 180 }\] you multiply it by pi/180

OpenStudy (anonymous):

okay..so my answer should be 4.712?

OpenStudy (johnweldon1993):

Not quite...you try to keep in in a fraction form...but simplify it down \[\frac{ 270\pi }{ 180 }\] can that be simplified at all?

OpenStudy (anonymous):

umm, how?

OpenStudy (johnweldon1993):

*hint....looks like both numbers can be divide by 90 right? so how many times does 90 go into 270? and how many times does 90 go into 180?

OpenStudy (anonymous):

3 and 2

OpenStudy (johnweldon1993):

right....so simplified down we have \[\frac{ 3\pi }{ 2 }\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

that makes a lot of sense.. thank you for your help

OpenStudy (johnweldon1993):

I hope it did! you're welcome